### v1

# ChartMaster: Advancing Chart-to-Code Generation with Real-World Charts and Chart Similarity Reinforcement Learning

Abstract
The chart-to-code generation task requires MLLMs to convert chart images into executable code. This task faces two major challenges: limited data diversity and insufficient maintenance of visual consistency between generated and original charts during training. Existing datasets mainly rely on seed data to prompt GPT models for code generation, resulting in homogeneous samples. To address this, we propose ReChartPrompt, which leverages real-world, human-designed charts from arXiv papers as prompts instead of synthetic seeds. Using the diverse styles and rich content of arXiv charts, we construct ReChartPrompt-240K, a large-scale and highly diverse dataset. Another challenge is that although SFT effectively improve code understanding, it often fails to ensure that generated charts are visually consistent with the originals. To address this, we propose ChartSimRL, a GRPO-based reinforcement learning algorithm guided by a novel chart similarity reward. This reward consists of attribute similarity, which measures the overlap of chart attributes such as layout and color between the generated and original charts, and visual similarity, which assesses similarity in texture and other overall visual features using convolutional neural networks. Unlike traditional text-based rewards such as accuracy or format rewards, our reward considers the multimodal nature of the chart-to-code task and effectively enhances the model’s ability to accurately reproduce charts. By integrating ReChartPrompt and ChartSimRL, we develop the ChartMaster model, which achieves state-of-the-art results among 7B-parameter models and even rivals GPT-4o on various chart-to-code generation benchmarks. All resources are available at https://github.com/WentaoTan/ChartMaster.
Figure 1: Fine-tuning Qwen2.5-VL-7B with ReChartPrompt data significantly enhances performance, and applying ChartSimRL further elevates it to match or exceed GPT-4o on certain metrics. For better representation, the “Rating” metric in the Plot2Code benchmark is multiplied by 10.
1 Introduction
The chart-to-code generation task aims to automatically convert chart images into executable code yang2024chartmimic, enabling applications including automated data analysis, report generation, and intelligent question answering zhao2025chartcoder. This task is challenging as it requires accurate visual understanding, cross-modal reasoning, and advanced code synthesis. Although recent advances in Multimodal Large Language Models (MLLMs) show promising results in various vision-language tasks, their performance on chart-to-code generation remains limited due to the unique complexity of charts and the need for precise code output.
Prior work such as ChartCoder zhao2025chartcoder advanced the field by building the large Chart2Code-160K dataset. This dataset is synthesized by guiding GPT-4o hurst2024gpt with predefined chart attributes like chart type, color, and text. While this approach reduces reliance on costly manual annotations and achieves strong performance, relying on predefined attribute seeds can introduce homogeneity and limit variability in the resulting dataset (see Appendix Fig. 5), potentially restricting model generalization to diverse real-world charts.
To address this limitation, we introduce Real-world Chart Prompt Code Generation (ReChartPrompt), a novel automated pipeline that extracts real-world chart images from arXiv papers and uses the Qwen2.5-VL-72B model bai2025qwen2 to generate code. Specifically, we collect 30,071 papers and use their charts as prompts to build a large-scale ReChartPrompt-240K dataset. As the charts in each paper are manually designed by their authors and cover a wide range of research fields, the dataset features diverse expertise and design styles. This rich diversity enables effective generalization to real-world scenarios. As shown in Fig. 1, fine-tuning with ReChartPrompt-240K improves performance on the ChartMimic benchmark by about 30 to 40 points.
While supervised fine-tuning with diverse data can help models generate better chart code, such token-level optimization cannot guarantee that the output charts are visually identical to the references. As shown in Fig. 4, SFT model produces charts closer to the ground truth than the baseline, but differences remain. To address this, we propose ChartSimRL, a reinforcement learning algorithm based on Group Relative Policy Optimization (GRPO) shao2024deepseekmath with a novel chart similarity reward. Specifically, the reward jointly considers attribute similarity—which evaluates the consistency of chart elements such as textual content, numerical values, layout, and color—and visual similarity, which assesses overall visual resemblance using convolutional neural networks (e.g., ResNet he2016deep) to extract and compare visual features. To the best of our knowledge, this is the first reward system designed to effectively capture the multimodal visual-semantic consistency that is crucial for chart-to-code generation. This enables models to produce charts that are not only code-accurate but also faithful to the originals, addressing a critical gap in prior research and supporting more robust generalization to real-world chart reproduction scenarios. As shown in Fig. 1, applying ChartSimRL to SFT-trained models brings performance close to that of GPT-4o.
In summary, we present ChartMaster, an efficient chart-to-code generation framework integrating the ReChartPrompt data generation pipeline and ChartSimRL reinforcement learning strategy. Built on the Qwen2.5-VL-7B model, ChartMaster achieves near GPT-4o performance on multiple benchmarks, advancing automation in chart reproduction. Our main contributions are: (1) ReChartPrompt, an automated data generation method leveraging real academic charts for diverse datasets; (2) ChartSimRL, a reinforcement learning algorithm combining visual and attribute similarity rewards; and (3) ChartMaster-7B, a compact model achieving near GPT-4o performance with only 7B parameters.
2 Related Work
2.1 Multimodal Code Generation
Recently, MLLMs have attracted increasing attention for their code generation capabilities zhang2024humaneval. For example, MMCode li2024mmcode focuses on algorithmic problems in visually rich contexts, where each task is accompanied by one or more images.
Among various code generation tasks, chart-to-code generation has emerged as a critical research area zhao2025chartcoder; yang2024matplotagent. Relevant benchmarks include Design2Code si2024design2code, which evaluates HTML webpage generation using CLIP scores and HTML block comparison metrics, and Plot2Code wu2024plot2code, which assesses both code correctness and visual similarity for comprehensive evaluation. However, since the datasets for Design2Code and Plot2Code are sourced from the web, there is a risk of data leakage, which may compromise the reliability of model evaluation. To address this issue, ChartMimic yang2024chartmimic provides a manually curated dataset of 4,800 chart-code pairs, along with additional fine-grained evaluation metrics.
Despite these benchmarks, large-scale chart-to-code training datasets remain scarce. ChartCoder zhao2025chartcoder addresses this by creating Chart2Code-160K, the first large-scale training set generated by guiding GPT-4o with predefined chart attributes such as type, color, values, and titles. It also uses the “Snippet of Thought” method zheng2023outline; luo2024python to break down code generation into finer steps, significantly enhancing the model’s chart reasoning ability. Nevertheless, the limited diversity of predefined attributes results in relatively homogeneous charts. To overcome this limitation, we propose ReChartPrompt, which leverages real charts from arXiv papers instead of synthetic seeds to generate more diverse and representative chart-code pairs.
2.2 Reinforcement Learning for MLLMs
Reinforcement learning (RL) has been proven effective in enhancing model capabilities wang2024comprehensive; milani2024explainable. For example, reinforcement learning from human feedback (RLHF) bai2022training and direct preference optimization (DPO) rafailov2023direct align model outputs with human preferences, significantly improving complex reasoning ability and output quality.
Building on these advances, Group Relative Policy Optimization (GRPO) shao2024deepseekmath was developed as a novel RL algorithm that updates policies based on relative rewards computed from groups of samples. For instance, DeepSeek-R1 guo2025deepseek uses simple but effective reward functions based on output accuracy and response format, enabling stable training and fostering emergent reasoning skills such as reflection and “a-ha” moments.
Inspired by DeepSeek-R1’s success, recent work has extended GRPO-based RL to MLLMs tan2025reason; zhang2025r1; peng2025lmm; shen2025vlm. These efforts can be categorized into two groups. The first adapts R1’s approach to MLLMs; for example, Vision-R1 huang2025vision constructs supervised data with reflection for cold-start training, then applies GRPO with accuracy- and format-based rewards. Similarly, MM-EUREKA meng2025mm improves reward design and loss functions, successfully reproducing the visual “aha moment” where the model revisits images and remarks “upon closer inspection.” While impactful, these works mainly target mathematical reasoning tasks. The second category applies GRPO to a wider range of tasks, including visual perception yu2025perception, segmentation liu2025seg, and grounding zhang2025improving, demonstrating GRPO’s robustness and generalizability across diverse domains.
However, to our knowledge, no prior work has applied GRPO to chart-to-code generation. The core difficulty lies in designing an effective reward function that encourages generated code to faithfully reproduce the given chart, both semantically and visually. To this end, we design a novel chart similarity reward that measures the alignment between generated and reference charts, leading to significant improvements in chart reproduction quality.
3 Method
Fig. 2 illustrates the overall framework of ChartMaster, which consists of two main stages: data generation and model training.
Figure 2: The overall framework of ChartMaster. (a) Real-world charts from arXiv are curated to create the ReChartPrompt-240K dataset for SFT (SFT is omitted in the figure). (b) The model is further optimized with ChartSimRL. (c) The definition of Chart Similarity; $\mathcal{G}$ denotes the semantic attribute extraction tool; $\mathcal{F}$ is the CNN-based feature extractor; and $\mathbf{f}$ is the extracted feature vector.
3.1 Using Real-World Charts to Generate Dataset
To improve dataset diversity, we use real-world chart images as input to guide code generation. This approach captures richer styles and content that predefined attribute seeds cannot represent.
(1) Collecting Images from arXiv. We leverage the arXiv API and Python’s requests library to download paper source files, including LaTeX sources and image files (.pdf, .png, .jpg). To ensure diversity, we query source files related to top conferences (e.g., ICLR) and journals (e.g., TPAMI), extracting all images for subsequent processing.
(2) Filtering Non-Chart Images. Since extracted images include various diagrams beyond charts, we use the Qwen2.5-VL-72B model to classify images into 12 predefined chart categories. Images outside these categories are discarded. Classification is performed by prompting the model with $P_{\text{type}}$ (see Fig. 6 in Appendix) to assign chart types.
(3) Generating code with ReChartPrompt. The Qwen2.5-VL-72B model has demonstrated strong chart-to-code generation capabilities. As an open-source model, it is easily deployed via the vLLM framework kwon2023efficient, making it well-suited for large-scale data generation. We design a set of 20 chart-to-code prompts to enrich instruction diversity, collectively referred to as $\mathbf{P}_{\text{rechart}}$ (see Fig. 7 in Appendix). Below is an example: <<Real-World Chart>>Please generate Python matplotlib code to recreate the picture shown.
(4) Code Execution, Filtering, and Dataset Construction. Generated code snippets may suffer from two issues: (a) execution errors caused by non-existent packages or syntax mistakes, and (b) discrepancies between the generated charts and the original images. To mitigate these problems, we execute all generated code and discard those that fail at runtime. We then pair the successfully executed code outputs with their generated images and instructions to form the final training triplets.
Summary. We download 30,071 papers from arXiv, extracting 288,992 chart images. The Qwen2.5-VL-72B model generates code for these charts, and after executing and filtering out failed cases, we obtain 242,479 high-quality triplets, which form the ReChartPrompt-240K dataset. Formally, the dataset is represented as $\mathcal{D}=\{(I_{i},T_{i},Y_{i})\}_{i=1}^{N}$, where $I_{i}$ denotes the chart image, $T_{i}\in\mathbf{P}_{\text{rechart}}$ is the instruction prompt, and $Y_{i}$ is the executable code.
3.2 Training ChartMaster: SFT and ChartSimRL
ChartMaster is trained in two stages: (1) SFT on the ReChartPrompt-240K dataset to establish a solid foundation; and (2) further optimized with ChartSimRL to address the limitations of SFT’s next-token prediction in maintaining visual consistency.
Supervised Fine-Tuning. We conduct SFT by maximizing the likelihood of ground-truth code $Y_{i}$ given chart image $I_{i}$ and instruction $T_{i}$:
$J_{\mathrm{SFT}}(\theta)=-\frac{1}{N}\sum_{i=1}^{N}\log\pi_{\theta}(Y_{i}\mid I_{i},T_{i}).$
Reinforcement Learning with ChartSimRL. While SFT strengthens the model’s basic capability, discrepancies may still exist between the generated charts and the originals (see Fig. 4). To further improve reproduction fidelity, we continue training the model using ChartSimRL. Specifically, for each training sample $(I,T,Y)$, the model samples a group of $M$ candidate codes:
$\{O_{1},O_{2},\dots,O_{M}\}\sim\pi_{\theta}(\cdot\mid I,T).$
Each candidate code $O_{i}$ is then executed to generate a chart image $\hat{I}_{i}$, which is subsequently compared with the original chart ${I}_{i}$ to compute a Chart Similarity Reward. If the execution of $O_{i}$ fails, the corresponding reward is set to zero.
Chart Similarity Reward. Traditional reward functions, such as the accuracy reward used in guo2025deepseek; huang2025vision, primarily assess the consistency between generated text and ground-truth text. However, the chart-to-code task is inherently multimodal, involving both code and generated charts, requiring evaluation of not only semantic correctness but also visual alignment. To this end, we design a novel chart similarity reward as:
$R_{i}=R_{i}^{\mathrm{attr}}+R_{i}^{\mathrm{vis}}.$
Here, $R_{i}^{\mathrm{attr}}$ measures the semantic consistency, and $R_{i}^{\mathrm{vis}}$ captures visual similarity.
Attribute Similarity: We develop a semantic attribute extraction tool, denoted $\mathcal{G}(\cdot)$, to obtain attribute sets from chart images and their code. Given $\mathcal{A}_{i}=\mathcal{G}(\hat{I}_{i},O_{i})$ and $\mathcal{A}^{*}=\mathcal{G}(I,Y)$, the semantic similarity $R_{i}^{\mathrm{attr}}$ is computed as their Jaccard similarity:
$R_{i}^{\mathrm{attr}}=\frac{|\mathcal{A}_{i}\cap\mathcal{A}^{*}|}{|\mathcal{A}_{i}\cup\mathcal{A}^{*}|}\in[0,1].$
By design, $R_{i}^{\mathrm{attr}}=1$ indicates a perfect match of semantic attributes, while lower values reflect semantic discrepancies. To accommodate minor numerical variations, we consider numerical values $a\in\mathcal{A}_{i}$ and $b\in\mathcal{A}^{*}$ matching if $|a-b|\leq 0.01\times|b|$.
Visual Similarity: We use a pretrained ResNet-18 network he2016deep $\mathcal{F}=\{\mathcal{F}_{1},\mathcal{F}_{2},\mathcal{F}_{3},\mathcal{F}_{4}\}$ to extract feature maps from both $I$ and $\hat{I}_{i}$. Here, $\mathcal{F}_{k}(\cdot)\in\mathbb{R}^{C_{k}\times H_{k}\times W_{k}}$ denotes the output feature map of the $k$-th residual block. We extract the feature map and flatten them into vectors like:
$F_{k}=\mathcal{F}_{k}(I),\quad\quad$ $\hat{F}_{k}^{(i)}=\mathcal{F}_{k}(\hat{I}_{i}),$
$\mathbf{f}_{k}=\mathrm{vec}(F_{k})\in\mathbb{R}^{d_{k}},\quad\quad$ $\hat{\mathbf{f}}_{k}^{(i)}=\mathrm{vec}(\hat{F}_{k}^{(i)})\in\mathbb{R}^{d_{k}},$
where $d_{k}=C_{k}\times H_{k}\times W_{k}$. The visual similarity reward is defined as the average cosine similarity between the corresponding feature vectors,
$R_{i}^{\mathrm{vis}}=\frac{1}{4}\sum_{k=1}^{4}\frac{\mathbf{f}_{k}\cdot\hat{\mathbf{f}}_{k}^{(i)}}{\|\mathbf{f}_{k}\|\,\|\hat{\mathbf{f}}_{k}^{(i)}\|}\in[0,1].$
Chart Similarity Reinforcement Learning. We normalize rewards within a group of $M$ candidates to compute relative advantages:
$\hat{A}_{i}=\frac{R_{i}-\mathrm{mean}(\{R_{j}\}_{j=1}^{M})}{\mathrm{std}(\{R_{j}\}_{j=1}^{M})},$
where $\mathrm{mean}(\cdot)$ and $\mathrm{std}(\cdot)$ denote the sample mean and standard deviation, respectively.
Following the GRPO framework shao2024deepseekmath, we update the model by maximizing the clipped surrogate objective with a KL penalty to stabilize training:
$J_{\mathrm{ChartSimRL}}(\theta)$ $=\mathbb{E}_{(I,T)\sim p_{\mathcal{D}},\,\{o_{i}\}_{i=1}^{M}\sim\pi_{\mathrm{old}}(\cdot|I,T)}\Bigg{[}\frac{1}{M}\sum_{i=1}^{M}\min\Big{(}\frac{\pi_{\theta}(o_{i}|I,T)}{\pi_{\mathrm{old}}(o_{i}|I,T)}\hat{A}_{i},$
$\operatorname{clip}\Big{(}\frac{\pi_{\theta}(o_{i}|I,T)}{\pi_{\mathrm{old}}(o_{i}|I,T)},1-\epsilon,1+\epsilon\Big{)}\hat{A}_{i}\Big{)}-\beta D_{\mathrm{KL}}\big{(}\pi_{\theta}(\cdot|I,T)\|\pi_{\mathrm{ref}}(\cdot|I,T)\big{)}\Bigg{]},$
where $\pi_{\mathrm{old}}$ is the previous policy, $\pi_{\mathrm{ref}}$ is the reference policy, $\epsilon$ is the clipping hyperparameter, and $\beta$ controls the KL regularization strength.
ChartSimRL guides the model to generate chart code that better aligns with the original charts’ semantic and visual properties, significantly improving chart-to-code generation performance beyond what is achievable by supervised fine-tuning alone.
4 Experiment
4.1 Implementation Details
During the collection of arXiv papers, we explicitly exclude papers that are used as benchmarks to avoid potential data leakage. In the data generation stage, we apply a greedy sampling strategy to filter chart data, retaining only images of 12 predefined chart types and discarding all others. Then, we randomly select an instruction from $\mathbf{P}_{\text{rechart}}$ to prompt the Qwen2.5-VL-72B bai2025qwen2 model to generate code via nucleus sampling, with a temperature of 0.1 and a top-p of 0.9.
For training, we use the Qwen2.5-VL-7B model bai2025qwen2 in two stages. In Stage 1, we perform SFT on the entire ReChartPrompt-240K dataset with a learning rate of $2\times 10^{-5}$, batch size 128, and a cosine annealing scheduler for one epoch; the resulting model is saved for Stage 2. In Stage 2, ChartSimRL training is conducted on 10% of the ReChartPrompt-240K dataset, using a smaller learning rate of $5\times 10^{-6}$ and generating $M=4$ candidate codes per sample. Candidate sampling uses temperature 1.0, top-p 1.0, and top-k 80 to encourage diversity. The batch size remains 128 (32 samples $\times$ 4 candidates each).
4.2 Evalution Benchmarks
Following zhao2025chartcoder, we assess the model’s chart-to-code generation performance on multiple benchmarks. ChartMimic Direct Mimic Task yang2024chartmimic: This benchmark includes 600 chart images. GPT-4o scores (0–100) serve as high-level similarity metrics. Additionally, low-level F1 scores for text, layout, chart type, and color are computed from code execution for fine-grained analysis. Plot2Code Direct Asking wu2024plot2code: Metrics include code pass rate, text match rate, and a 10-point GPT-4V visual similarity score, jointly assessing code correctness and visual fidelity. ChartX Chart Redrawing Task xia2024chartx: This benchmark uses GPT-4 (0–5 scale) to evaluate code-generated chart redrawings.
Table 1: Evaluation results of various MLLMs. Reported results are taken from existing benchmarks when available; missing results are supplemented by our own experiments. Among open-source 7B-scale models, our method achieves the best performance.
Model ChartMimic Plot2Code ChartX
Exec.Rate Low-Level High-Level Pass Rate Text-Match Rating GPT-score
Full score 100 100 100 100 100 10 5
Closed-Source Model
GeminiProVision team2023gemini 68.2 53.8 53.3 68.2 53.6 3.69 -
Claude-3-opus anthropic2024claude 83.3 60.5 60.1 84.1 57.5 3.80 -
GPT-4V hurst2024gpt 91.2 76.4 78.9 84.1 57.7 5.58 2.63
GPT-4o hurst2024gpt 93.2 79.0 83.5 88.6 56.3 5.71 2.36
Open-Source Model
ChartAssisstant-13B meng2024chartassisstant - - - - - - 0.82
ChartVLM-L-14B xia2024chartx 19.5 15.8 13.9 - - - 1.58
DeepSeek-VL-7B lu2024deepseek 41.3 19.0 17.6 64.4 32.6 2.26 -
TinyChart-3B zhang2024tinychart 42.5 26.3 25.9 43.2 44.6 2.19 1.89
ChartLlama-13B han2023chartllama 57.5 24.8 28.1 58.4 40.3 2.32 0.94
LLaVA-Next-Mistral-7B li2024llava 59.7 20.7 21.3 72.0 38.7 2.87 -
InternVL2-8B chen2024internvl 61.8 34.4 38.9 77.3 37.1 2.78 1.63
Qwen2-VL-7B wang2024qwen2 67.0 32.9 35.0 68.2 33.8 3.10 1.50
MiniCPM-Llama3-V2.5-8B yao2024minicpm 80.3 36.6 42.1 76.3 37.3 2.61 1.66
Qwen2-VL-72B wang2024qwen2 73.3 54.4 50.9 72.0 53.4 4.26 1.69
InternVL2-Llama3-76B chen2024internvl 83.2 54.8 62.2 85.6 46.6 3.89 1.74
Qwen2.5-VL-72B bai2025qwen2 88.5 72.7 71.8 84.8 68.4 6.83 2.52
ChartCoder-7B zhao2025chartcoder 91.4 77.4 74.0 87.9 54.5 4.50 2.09
Qwen2.5-VL-7B (Baseline) bai2025qwen2 65.5 39.9 34.2 67.4 43.8 4.60 2.18
ChartMaster-7B 93.8 78.2 77.3 88.2 62.6 5.65 2.46
4.3 Comparison with SOTA
Table 1 compares our ChartMaster-7B with various MLLMs on multiple chart-to-code benchmarks. ChartMaster-7B achieves SOTA results among open-source 7B-scale models, even approaching the performance of GPT-4o. Notably, ChartMaster-7B consistently outperforms the baseline Qwen2.5-VL-7B across all metrics; for example, on the ChartMimic benchmark, it improves both low-level and high-level metrics by nearly 40 percentage points, highlighting the effectiveness of our approach. Furthermore, although our training dataset is sourced from the much larger Qwen2.5-VL-72B model—essentially a distillation-like setting—ChartMaster-7B still surpasses Qwen2.5-VL-72B on several benchmarks. These results demonstrate the strength of the ChartMaster framework in narrowing, and in some cases closing, the performance gap with significantly larger models.
Table 2: Ablation study on the contribution of each key component.
ReChartPrompt ChartSimRL ChartMimic Plot2Code ChartX
Exec.Rate Low-Level High-Level Pass Rate Text-Match Rating GPT-score
65.5 39.9 34.2 67.4 43.8 4.60 2.18
$\checkmark$ 91.1 73.7 73.3 80.3 59.3 5.34 2.36
$\checkmark$ $\checkmark$ 93.8 78.2 77.3 88.2 62.6 5.65 2.46
4.4 Ablation Study
Ablation study on ChartMaster. To assess the contribution of each component, we conduct an ablation study as summarized in Table 2. The base Qwen2.5-VL-7B model, without ReChartPrompt or ChartSimRL, demonstrates limited performance across benchmarks, revealing its restricted ability in both code generation and visual/semantic understanding. Introducing the ReChartPrompt-240K dataset leads to significant improvements in all metrics, demonstrating the high quality and effectiveness of ReChartPrompt-240K for chart-to-code generation. Further applying ChartSimRL on top of ReChartPrompt brings consistent gains, especially in GPT-based evaluation scores. These results demonstrate that ChartSimRL effectively guides the model to produce code that better matches the original charts in both semantic and visual aspects.
Table 3: Ablation study of the Attribute and Visual similarity components in ChartSimRL.
$R_{i}^{\mathrm{attr}}$ $R_{i}^{\mathrm{vis}}$ ChartMimic
Exec.Rate Low-Level High-Level
91.1 73.7 73.3
$\checkmark$ 92.1 76.2 75.2
$\checkmark$ 92.1 77.7 76.0
$\checkmark$ $\checkmark$ 93.8 78.2 77.3
Table 4: Ablation study of different attribute similarity metrics on the ChartMimic benchmark.
$R_{i}^{\mathrm{attr}}$ Formula ChartMimic
Exec.Rate Low-Level High-Level
Baseline - 91.1 73.7 73.3
Precision $\frac{|\mathcal{A}_{i}\cap\mathcal{A}^{*}|}{|\mathcal{A}_{i}|}$ 90.0 72.6 72.9
Recall $\frac{|\mathcal{A}_{i}\cap\mathcal{A}^{*}|}{|\mathcal{A}^{*}|}$ 90.6 74.7 74.8
F1 $\frac{|\mathcal{A}_{i}\cap\mathcal{A}^{*}|}{(|\mathcal{A}_{i}|+|\mathcal{A}^{*}|)/2}$ 91.6 75.4 75.3
Jaccard $\frac{|\mathcal{A}_{i}\cap\mathcal{A}^{*}|}{|\mathcal{A}_{i}\cup\mathcal{A}^{*}|}$ 92.1 76.2 75.2
Ablation study on ChartSimRL. ChartSimRL introduces a novel multimodal chart similarity reward, which goes beyond conventional text-based rewards by integrating both semantic and visual similarity between the candidate and original charts. To analyze the contribution of each component, we conduct ablation experiments as shown in Table 4. The results indicate that using either $R_{i}^{\mathrm{attr}}$ or $R_{i}^{\mathrm{vis}}$ alone improves all performance. Among them, the visual similarity reward leads to more pronounced gains, highlights the critical role of visual fidelity in chart-to-code generation. Furthermore, combining both rewards yields the best results, demonstrating the effectiveness of a multi-faceted reward design.
Ablation study on Attribute Similarity. We further evaluate the impact of different attribute similarity metrics—Precision, Recall, F1 score, and Jaccard similarity—on the ChartMimic benchmark (Table 4). Results show that optimizing solely for Precision can slightly decrease certain metrics, indicating that focusing only on attribute correctness without considering coverage may be suboptimal. In contrast, Recall, F1, and Jaccard similarity all lead to better performance, with Jaccard similarity achieving the highest overall scores.
Notably, while the F1 score balances Precision and Recall through the harmonic mean, Jaccard similarity measures the intersection over union between predicted and reference attribute sets, imposing a stricter penalty on missing or redundant attributes. This stricter overlap requirement allows Jaccard similarity to better capture the true semantic similarity between attribute sets, thus more effectively guiding the model’s learning process.
Ablation Study on Visual Similarity. Table 5 presents an ablation study of various visual similarity metrics on the ChartMimic benchmark. Standard metrics like MSE, SSIM wang2004image, and PSNR hore2010image, mainly assessing pixel-level or structural similarity, generally perform worse than advanced methods. In particular, SSIM underperforms significantly, indicating that pixel-based measures struggle to capture the complex visual nuances required for chart generation.
In contrast, CNN-based metrics like AlexNet krizhevsky2012imagenet, VGG simonyan2014very and ResNet he2016deep that compare features in a learned representation space consistently outperform the baseline and pixel-level metrics across all evaluation criteria. ResNet-18 achieves the highest performance, highlighting the effectiveness of deep visual features.
Table 5: Ablation study of different visual similarity metrics on the ChartMimic benchmark.
$R_{i}^{\mathrm{vis}}$ ChartMimic
Exec.Rate Low-Level High-Level
Baseline 91.1 73.7 73.3
Standard Metrics:
MSE 91.1 73.6 72.5
SSIM 82.5 65.2 67.0
PSNR 91.4 75.1 73.4
CNN-Based Metrics:
AlexNet 90.3 74.7 74.1
VGG 91.3 75.5 74.9
ResNet-18 92.1 77.7 76.0
MLLM-Based Metrics:
Qwen-2.5-VL-72B 91.7 77.5 74.7
For MLLM-based metrics, we use high-level similarity prompts from ChartMimic with Qwen-2.5-VL-72B to evaluate the similarity between generated and reference charts. These metrics further improve upon the baseline, especially in Execution Rate and Low-Level similarity. However, they still fall short of the best CNN-based metrics in High-Level similarity, suggesting that while MLLMs have strong semantic understanding, further optimization is needed for specialized visual tasks.
Figure 3: Comparison of dataset diversity and model performance of Qwen2.5-VL-7B after fine-tuning with Chart2Code-160K and ReChartPromp-160K datasets.
Comparison with Advanced Dataset. To thoroughly evaluate the diversity and quality of our dataset, we compare it with Chart2Code-160K. For fairness, 160K samples are randomly selected from our data to form the ReChartPromp-160K subset. Results are shown in Fig. 3. The left subfigure displays normalized counts of unique attributes—Numeric, Color, Text, and Layout—extracted from both datasets using tool $\mathcal{G}(\cdot)$. Higher counts indicate greater attribute diversity. ReChartPromp-160K exhibits substantially more unique attributes across all categories, especially in Text and Layout. While Chart2Code-160K relies on seed data, often resulting in repetitive attributes (see Appendix Fig. 5), ReChartPromp-160K draws from 30,071 diverse arXiv papers, yielding richer and less redundant attribute distributions. This greater diversity leads to tangible benefits: as shown in the right subfigure, models trained on ReChartPromp-160K consistently outperform those trained on Chart2Code-160K, demonstrating the importance of attribute diversity for robust and effective chart understanding and code generation.
4.5 Qualitative analysis
Figure 4: The test results of various models on the ChartMimic benchmark. “Base.+ReCha.” refers to the baseline model fine-tuned with the ReChartPrompt-240K dataset. Incorporating ReChartPrompt significantly enhances the chart-to-code generation capability of the base model, while ChartSimRL further improves the handling of fine details.
To further evaluate the effectiveness of our proposed approach, Fig. 4 presents qualitative comparisons of chart outputs generated by different models on the ChartMimic benchmark. The results reveal several key insights: (1) The baseline model can generate basic chart layouts, but often fails to reproduce fine-grained visual details. Consequently, its outputs frequently diverge from the reference charts. (2) ReChartPrompt-240K enhances chart-to-code generation accuracy. Fine-tuning the base model with our ReChartPrompt-240K dataset (“Base.+ReCha.”) makes the generated charts more accurate and diverse. This shows that using real-world chart-code pairs helps the model better adapt to different chart types and styles. However, there are still minor issues, like slight color or element position mismatches, indicating that standard supervised learning alone cannot achieve perfect visual consistence. (3) ChartSimRL further improves visual and semantic alignment. Incorporating ChartSimRL algorithm enables the model to more closely align its outputs with the ground truth. For example, the color usage in chart 1 and the arrow style in chart 2 better match the ground truth.
In summary, by combining ReChartPrompt and ChartSimRL, ChartMaster achieves visually accurate and semantically precise chart code, demonstrating strong generalization. This integrated approach also significantly narrows the gap with models like GPT-4o, highlighting ChartMaster’s practical effectiveness in chart-to-code generation.
5 Conclusion
In this paper, we propose ChartMaster, including a novel chart-to-code generation framework and a customized reinforcement learning algorithm. By introducing ReChartPrompt, we overcome the data homogeneity issues of previous works and construct a highly diverse ReChartPrompt-240K dataset. Our ChartSimRL algorithm integrates both semantic and visual similarity rewards, enabling the model to generate chart code that closely matches the original charts. Experimental results demonstrate that ChartMaster achieves performance competitive with GPT-4o on the chart-to-code generation task.
Beyond its technical contributions, ChartMaster can facilitate automated scientific reporting and support data-driven decision-making across a wide range of domains. While our current framework focuses on common chart types and Python-based code, extending its capabilities to support a broader variety of chart formats and programming languages remains an important direction for future work.
Appendix A Appendix
Figure 5: Dataset visualization. The charts in Chart2Code-160K exhibit homogenization, which affects diversity; the charts in ReChartPrompt demonstrate greater variety, especially in terms of textual content within the tables and layout attributes.
Figure 6: Prompt used for chart type classification ($P_{\text{type}}$). The Qwen2.5-VL-72B model is prompted with this template to assign each image to one of 12 predefined chart categories.
Figure 7: Prompt for chart-to-code generation ($\mathbf{P}_{\text{rechart}}$). Twenty diverse prompts are designed to instruct the Qwen2.5-VL-72B model to generate Python matplotlib code from chart images, enhancing instruction diversity.
Figure 8: More test results of various models on the ChartMimic benchmark.

### v2

@@
 The chart-to-code generation task requires MLLMs to convert chart images into executable code.
-This task faces two major challenges: limited data diversity and insufficient maintenance of visual consistency between generated and original charts during training.
-Existing datasets mainly rely on seed data to prompt GPT models for code generation, resulting in homogeneous samples.
-To address this, we propose ReChartPrompt, which leverages real-world, human-designed charts from arXiv papers as prompts instead of synthetic seeds.
-Using the diverse styles and rich content of arXiv charts, we construct ReChartPrompt-240K, a large-scale and highly diverse dataset.
-Another challenge is that although SFT effectively improve code understanding, it often fails to ensure that generated charts are visually consistent with the originals.
+This task faces two main challenges: limited data diversity and the difficulty of maintaining visual consistency between generated charts and the original ones.
+Existing datasets mainly rely on synthetic seed data to prompt GPT models for code generation, resulting in homogeneous samples that limit model generalization to real-world chart styles.
+To address this, we propose ReChartPrompt, leveraging real-world, human-designed charts extracted from arXiv papers as prompts.
+By harnessing the rich content and diverse visual styles of arXiv charts, we construct ReChartPrompt-240K, a large-scale and highly diverse dataset that better reflects realistic chart variations.
+For the second challenge, although SFT improves code understanding by optimizing next-token prediction, it does not provide direct supervision on visual features.
+As a result, it often fails to guarantee that the generated charts visually match the original ones.
 To address this, we propose ChartSimRL, a GRPO-based reinforcement learning algorithm guided by a novel chart similarity reward.
-This reward consists of attribute similarity, which measures the overlap of chart attributes such as layout and color between the generated and original charts, and visual similarity, which assesses similarity in texture and other overall visual features using convolutional neural networks.
-Unlike traditional text-based rewards such as accuracy or format rewards, our reward considers the multimodal nature of the chart-to-code task and effectively enhances the model’s ability to accurately reproduce charts.
-By integrating ReChartPrompt and ChartSimRL, we develop the ChartMaster model, which achieves state-of-the-art results among 7B-parameter models and even rivals GPT-4o on various chart-to-code generation benchmarks.
+This reward consists of two components: attribute similarity, which measures the overlap of chart attributes like layout and color between the generated and original charts, and visual similarity, which evaluates overall visual features, including texture, using convolutional neural networks.
+Unlike traditional text-based rewards, our reward accounts for the multimodal nature of the chart-to-code generation task, significantly enhancing the model’s ability to accurately reproduce charts.
+Integrating ReChartPrompt and ChartSimRL, we develop the ChartMaster model, achieving SOTA results among 7B-parameter models and rivaling GPT-4o on various chart-to-code benchmarks.
 All resources are available at https://github.com/WentaoTan/ChartMaster.
-Figure 1: Fine-tuning Qwen2.5-VL-7B with ReChartPrompt data significantly enhances performance, and applying ChartSimRL further elevates it to match or exceed GPT-4o on certain metrics.
-For better representation, the “Rating” metric in the Plot2Code benchmark is multiplied by 10.
 1 Introduction
-The chart-to-code generation task aims to automatically convert chart images into executable code yang2024chartmimic, enabling applications including automated data analysis, report generation, and intelligent question answering zhao2025chartcoder.
+The chart-to-code generation task aims to automatically convert chart images into executable code (Yang et al., 2024a), enabling applications including automated data analysis, report generation, and intelligent question answering (Zhao et al., 2025).
 This task is challenging as it requires accurate visual understanding, cross-modal reasoning, and advanced code synthesis.
 Although recent advances in Multimodal Large Language Models (MLLMs) show promising results in various vision-language tasks, their performance on chart-to-code generation remains limited due to the unique complexity of charts and the need for precise code output.
-Prior work such as ChartCoder zhao2025chartcoder advanced the field by building the large Chart2Code-160K dataset.
-This dataset is synthesized by guiding GPT-4o hurst2024gpt with predefined chart attributes like chart type, color, and text.
-While this approach reduces reliance on costly manual annotations and achieves strong performance, relying on predefined attribute seeds can introduce homogeneity and limit variability in the resulting dataset (see Appendix Fig.
+Prior work, such as ChartCoder (Zhao et al., 2025), advanced the field by building the large Chart2Code-160K dataset.
+This dataset is synthesized by guiding GPT-4o (Hurst et al., 2024) with predefined chart attributes like chart type, color, and text.
+While this approach reduces the need for costly manual annotations and achieves strong performance, relying on predefined attribute seeds can introduce homogeneity and limit variability in the resulting dataset (see Appendix Fig.
 5), potentially restricting model generalization to diverse real-world charts.
-To address this limitation, we introduce Real-world Chart Prompt Code Generation (ReChartPrompt), a novel automated pipeline that extracts real-world chart images from arXiv papers and uses the Qwen2.5-VL-72B model bai2025qwen2 to generate code.
-Specifically, we collect 30,071 papers and use their charts as prompts to build a large-scale ReChartPrompt-240K dataset.
-As the charts in each paper are manually designed by their authors and cover a wide range of research fields, the dataset features diverse expertise and design styles.
-This rich diversity enables effective generalization to real-world scenarios.
+Figure 1: Performance comparison on three benchmarks.
+Our method outperforms ChartCoder-7B (Zhao et al., 2025), and matches or exceeds GPT-4o on certain metrics.
+For better representation, the “Rating” metric in the Plot2Code benchmark is multiplied by 10.
+To address this limitation, we introduce Real-world Chart Prompt Code Generation (ReChartPrompt), a novel automated pipeline that extracts real chart images from arXiv papers and leverages the Qwen2.5-VL-72B model (Bai et al., 2025) to generate corresponding code.
+By collecting 30,071 papers and utilizing their author-designed charts as prompts, we construct ReChartPrompt-240K, a large-scale dataset comprising 240K chart–code pairs.
+Since these charts originate from papers across diverse research fields and exhibit a wide variety of design styles, the dataset captures rich visual and semantic diversity, as illustrated in Fig.
+5.
+This heterogeneity enables effective generalization to real-world scenarios.
+While supervised fine-tuning with diverse data can help models generate better chart code, such next-token prediction alone does not ensure the output charts are visually faithful to the references.
 As shown in Fig.
-1, fine-tuning with ReChartPrompt-240K improves performance on the ChartMimic benchmark by about 30 to 40 points.
-While supervised fine-tuning with diverse data can help models generate better chart code, such token-level optimization cannot guarantee that the output charts are visually identical to the references.
-As shown in Fig.
-4, SFT model produces charts closer to the ground truth than the baseline, but differences remain.
-To address this, we propose ChartSimRL, a reinforcement learning algorithm based on Group Relative Policy Optimization (GRPO) shao2024deepseekmath with a novel chart similarity reward.
-Specifically, the reward jointly considers attribute similarity—which evaluates the consistency of chart elements such as textual content, numerical values, layout, and color—and visual similarity, which assesses overall visual resemblance using convolutional neural networks (e.g., ResNet he2016deep) to extract and compare visual features.
-To the best of our knowledge, this is the first reward system designed to effectively capture the multimodal visual-semantic consistency that is crucial for chart-to-code generation.
-This enables models to produce charts that are not only code-accurate but also faithful to the originals, addressing a critical gap in prior research and supporting more robust generalization to real-world chart reproduction scenarios.
-As shown in Fig.
-1, applying ChartSimRL to SFT-trained models brings performance close to that of GPT-4o.
-In summary, we present ChartMaster, an efficient chart-to-code generation framework integrating the ReChartPrompt data generation pipeline and ChartSimRL reinforcement learning strategy.
-Built on the Qwen2.5-VL-7B model, ChartMaster achieves near GPT-4o performance on multiple benchmarks, advancing automation in chart reproduction.
-Our main contributions are: (1) ReChartPrompt, an automated data generation method leveraging real academic charts for diverse datasets; (2) ChartSimRL, a reinforcement learning algorithm combining visual and attribute similarity rewards; and (3) ChartMaster-7B, a compact model achieving near GPT-4o performance with only 7B parameters.
+4, the SFT model produces charts closer to the ground truth than the baseline, but noticeable discrepancies remain in color, element positioning, and other visual attributes.
+To address this, we propose ChartSimRL, a reinforcement learning algorithm based on Group Relative Policy Optimization (GRPO) (Shao et al., 2024), guided by a novel chart similarity reward.
+Specifically, the reward jointly considers (1) attribute similarity that evaluates the consistency of chart elements such as textual content, numerical values, layout and color, and (2) visual similarity, which assesses holistic visual resemblance using convolutional neural networks (e.g., ResNet (He et al., 2016)) to extract and compare visual features.
+To the best of our knowledge, this is the first reward system that explicitly enforces multimodal visual-semantic consistency for chart-to-code generation.
+By encouraging models to produce code that renders charts both semantically accurate and visually faithful, we address a critical gap in prior research and support more robust generalization to real-world chart reproduction.
+In summary, we introduce ChartMaster, an efficient framework for chart-to-code generation that combines the ReChartPrompt data generation pipeline with the ChartSimRL reinforcement learning strategy.
+Our key contributions are: (1) ReChartPrompt, an automated method for generating diverse datasets from real-world charts; (2) ChartSimRL, a reinforcement learning algorithm that uses both visual and attribute similarity to improve output; and (3) ChartMaster-7B, a compact model that delivers near GPT-4o performance with only 7 billion parameters.
+Fig.
+1 highlights its efficiency and effectiveness.
 2 Related Work
 2.1 Multimodal Code Generation
-Recently, MLLMs have attracted increasing attention for their code generation capabilities zhang2024humaneval.
-For example, MMCode li2024mmcode focuses on algorithmic problems in visually rich contexts, where each task is accompanied by one or more images.
-Among various code generation tasks, chart-to-code generation has emerged as a critical research area zhao2025chartcoder; yang2024matplotagent.
-Relevant benchmarks include Design2Code si2024design2code, which evaluates HTML webpage generation using CLIP scores and HTML block comparison metrics, and Plot2Code wu2024plot2code, which assesses both code correctness and visual similarity for comprehensive evaluation.
+Multimodal large language models (MLLMs) have recently demonstrated strong capabilities in code generation (Zhang et al., 2024a; Tan et al., 2025b).
+Notably, MMCode (Li et al., 2024b) targets algorithmic problems embedded in visually rich contexts, where tasks are accompanied by one or more images.
+Among multimodal code generation tasks, chart-to-code translation has emerged as a critical challenge (Yang et al., 2024b).
+Existing benchmarks include Design2Code (Si et al., 2024), which evaluates HTML generation using CLIP scores (Radford et al., 2021) and structural HTML metrics, and Plot2Code (Wu et al., 2024), which assesses both code correctness and visual fidelity.
 However, since the datasets for Design2Code and Plot2Code are sourced from the web, there is a risk of data leakage, which may compromise the reliability of model evaluation.
-To address this issue, ChartMimic yang2024chartmimic provides a manually curated dataset of 4,800 chart-code pairs, along with additional fine-grained evaluation metrics.
+To address this issue, ChartMimic (Yang et al., 2024a) provides a manually curated dataset of 4,800 chart-code pairs, along with additional fine-grained evaluation metrics.
 Despite these benchmarks, large-scale chart-to-code training datasets remain scarce.
-ChartCoder zhao2025chartcoder addresses this by creating Chart2Code-160K, the first large-scale training set generated by guiding GPT-4o with predefined chart attributes such as type, color, values, and titles.
-It also uses the “Snippet of Thought” method zheng2023outline; luo2024python to break down code generation into finer steps, significantly enhancing the model’s chart reasoning ability.
-Nevertheless, the limited diversity of predefined attributes results in relatively homogeneous charts.
-To overcome this limitation, we propose ReChartPrompt, which leverages real charts from arXiv papers instead of synthetic seeds to generate more diverse and representative chart-code pairs.
+ChartCoder (Zhao et al., 2025) addresses this by creating Chart2Code-160K, the first large-scale training set generated by guiding GPT-4o with predefined chart attributes such as type, color, values, and titles.
+It further employs the “Snippet of Thought” strategy (Zheng et al., 2023; Luo et al., 2024) to decompose code generation into structured steps, significantly boosting chart reasoning.
+Yet, reliance on fixed attributes limits chart diversity.
+In contrast, our ReChartPrompt leverages real-world charts from arXiv papers as prompts, yielding more diverse and representative chart–code pairs.
 2.2 Reinforcement Learning for MLLMs
-Reinforcement learning (RL) has been proven effective in enhancing model capabilities wang2024comprehensive; milani2024explainable.
-For example, reinforcement learning from human feedback (RLHF) bai2022training and direct preference optimization (DPO) rafailov2023direct align model outputs with human preferences, significantly improving complex reasoning ability and output quality.
-Building on these advances, Group Relative Policy Optimization (GRPO) shao2024deepseekmath was developed as a novel RL algorithm that updates policies based on relative rewards computed from groups of samples.
-For instance, DeepSeek-R1 guo2025deepseek uses simple but effective reward functions based on output accuracy and response format, enabling stable training and fostering emergent reasoning skills such as reflection and “a-ha” moments.
-Inspired by DeepSeek-R1’s success, recent work has extended GRPO-based RL to MLLMs tan2025reason; zhang2025r1; peng2025lmm; shen2025vlm.
-These efforts can be categorized into two groups.
-The first adapts R1’s approach to MLLMs; for example, Vision-R1 huang2025vision constructs supervised data with reflection for cold-start training, then applies GRPO with accuracy- and format-based rewards.
-Similarly, MM-EUREKA meng2025mm improves reward design and loss functions, successfully reproducing the visual “aha moment” where the model revisits images and remarks “upon closer inspection.” While impactful, these works mainly target mathematical reasoning tasks.
-The second category applies GRPO to a wider range of tasks, including visual perception yu2025perception, segmentation liu2025seg, and grounding zhang2025improving, demonstrating GRPO’s robustness and generalizability across diverse domains.
-However, to our knowledge, no prior work has applied GRPO to chart-to-code generation.
-The core difficulty lies in designing an effective reward function that encourages generated code to faithfully reproduce the given chart, both semantically and visually.
-To this end, we design a novel chart similarity reward that measures the alignment between generated and reference charts, leading to significant improvements in chart reproduction quality.
+Reinforcement learning (RL) effectively enhances model capabilities Wang et al.
+(2024b); Milani et al.
+(2024); Tan et al.
+(2025c).
+For example, RL from human feedback (RLHF) Bai et al.
+(2022) and direct preference optimization (DPO) Rafailov et al.
+(2023) aligned model outputs with human preferences, improving complex reasoning and output quality.
+Building on these advances, Group Relative Policy Optimization (GRPO) Shao et al.
+(2024) was proposed as a novel RL algorithm that updated policies using relative rewards computed from groups of samples.
+DeepSeek-R1 Guo et al.
+(2025) employed simple yet effective rewards based on output accuracy and response format, which enabled stable training and emergent reasoning such as reflection and “a-ha” moments.
+Inspired by DeepSeek-R1’s success, recent work extended GRPO-based RL to MLLMs (Tan et al., 2025a; Zhang et al., 2025b; Peng et al., 2025; Shen et al., 2025) in two main directions.
+The first adapts R1’s method to MLLMs—for instance, Vision-R1 (Huang et al., 2025) uses SFT data with reflection for cold-start training and applies GRPO with accuracy- and format-based rewards.
+Similarly, MM-EUREKA (Meng et al., 2025) refines reward design and loss functions, successfully reproducing the visual “aha moment,” where the model revisits images “upon closer inspection.” These works primarily focus on mathematical reasoning tasks.
+The second direction applies GRPO to broader tasks such as visual perception (Yu et al., 2025), segmentation (Liu et al., 2025), and grounding (Zhang et al., 2025a), demonstrating its robustness and generalizability across domains.
+However, to our knowledge, GRPO has not been applied to chart-to-code generation, mainly due to the challenge of designing reward functions that encourage generated code to faithfully reproduce charts both semantically and visually.
+We address this by proposing a novel chart similarity reward, significantly improving chart reproduction quality.
 3 Method
@@
 (b) The model is further optimized with ChartSimRL.
-(c) The definition of Chart Similarity; $\mathcal{G}$ denotes the semantic attribute extraction tool; $\mathcal{F}$ is the CNN-based feature extractor; and $\mathbf{f}$ is the extracted feature vector.
+(c) The definition of Chart Similarity: $\mathcal{G}$ denotes the semantic attribute extraction tool; $\mathcal{F}$ is the CNN-based feature extractor; and $\mathbf{f}$ is the extracted feature vector.
 3.1 Using Real-World Charts to Generate Dataset
-To improve dataset diversity, we use real-world chart images as input to guide code generation. This approach captures richer styles and content that predefined attribute seeds cannot represent.
+To improve dataset diversity, we use real-world chart images as input to guide code generation, as shown in Fig.
+2 (a).
+This approach captures richer styles and content that predefined attribute seeds cannot represent.
 (1) Collecting Images from arXiv.
@@
 The Qwen2.5-VL-72B model has demonstrated strong chart-to-code generation capabilities.
-As an open-source model, it is easily deployed via the vLLM framework kwon2023efficient, making it well-suited for large-scale data generation.
+As an open-source model, it is easily deployed via the vLLM framework (Kwon et al., 2023), making it well-suited for large-scale data generation.
 We design a set of 20 chart-to-code prompts to enrich instruction diversity, collectively referred to as $\mathbf{P}_{\text{rechart}}$ (see Fig.
@@
 Summary.
-We download 30,071 papers from arXiv, extracting 288,992 chart images.
-The Qwen2.5-VL-72B model generates code for these charts, and after executing and filtering out failed cases, we obtain 242,479 high-quality triplets, which form the ReChartPrompt-240K dataset.
-Formally, the dataset is represented as $\mathcal{D}=\{(I_{i},T_{i},Y_{i})\}_{i=1}^{N}$, where $I_{i}$ denotes the chart image, $T_{i}\in\mathbf{P}_{\text{rechart}}$ is the instruction prompt, and $Y_{i}$ is the executable code.
+We download 30,071 papers from arXiv and extract their figures, filtering out non-chart ones to obtain 288,992 chart images.
+Using these charts, the Qwen2.5-VL-72B model generates corresponding code.
+After executing the generated code and removing failed cases, we collect 242,479 high-quality triplets that constitute the ReChartPrompt-240K dataset.
+Formally, the dataset is defined as $\mathcal{D}=\{(I_{i},T_{i},Y_{i})\}_{i=1}^{N}$, where $I_{i}$ represents a chart image, $T_{i}\in\mathbf{P}_{\text{rechart}}$ is the instruction prompt, and $Y_{i}$ denotes the executable code.
+Notably, all real-world chart data and generation models employed in this process are open-source, ensuring minimal cost and excellent scalability.
 3.2 Training ChartMaster: SFT and ChartSimRL
@@
 4).
-To further improve reproduction fidelity, we continue training the model using ChartSimRL.
+To further improve reproduction fidelity, we continue training the model using ChartSimRL, as illustrated in Fig.
+2 (b).
 Specifically, for each training sample $(I,T,Y)$, the model samples a group of $M$ candidate codes:
@@
 Chart Similarity Reward.
-Traditional reward functions, such as the accuracy reward used in guo2025deepseek; huang2025vision, primarily assess the consistency between generated text and ground-truth text.
+Traditional reward functions, such as the accuracy reward used in (Guo et al., 2025; Huang et al., 2025), primarily assess the consistency between generated text and ground-truth text.
 However, the chart-to-code task is inherently multimodal, involving both code and generated charts, requiring evaluation of not only semantic correctness but also visual alignment.
@@
 $R_{i}=R_{i}^{\mathrm{attr}}+R_{i}^{\mathrm{vis}}.$
-Here, $R_{i}^{\mathrm{attr}}$ measures the semantic consistency, and $R_{i}^{\mathrm{vis}}$ captures visual similarity.
+Here, $R_{i}^{\mathrm{attr}}$ measures the semantic consistency, and $R_{i}^{\mathrm{vis}}$ captures visual similarity (see Fig. 2 (c)).
 Attribute Similarity: We develop a semantic attribute extraction tool, denoted $\mathcal{G}(\cdot)$, to obtain attribute sets from chart images and their code.
@@
 To accommodate minor numerical variations, we consider numerical values $a\in\mathcal{A}_{i}$ and $b\in\mathcal{A}^{*}$ matching if $|a-b|\leq 0.01\times|b|$.
-Visual Similarity: We use a pretrained ResNet-18 network he2016deep $\mathcal{F}=\{\mathcal{F}_{1},\mathcal{F}_{2},\mathcal{F}_{3},\mathcal{F}_{4}\}$ to extract feature maps from both $I$ and $\hat{I}_{i}$.
+Visual Similarity: We use a pretrained ResNet-18 network (He et al., 2016) $\mathcal{F}=\{\mathcal{F}_{1},\mathcal{F}_{2},\mathcal{F}_{3},\mathcal{F}_{4}\}$ to extract feature maps from both $I$ and $\hat{I}_{i}$.
 Here, $\mathcal{F}_{k}(\cdot)\in\mathbb{R}^{C_{k}\times H_{k}\times W_{k}}$ denotes the output feature map of the $k$-th residual block.
@@
 where $\mathrm{mean}(\cdot)$ and $\mathrm{std}(\cdot)$ denote the sample mean and standard deviation, respectively.
-Following the GRPO framework shao2024deepseekmath, we update the model by maximizing the clipped surrogate objective with a KL penalty to stabilize training:
+Following the GRPO framework (Shao et al., 2024), we update the model by maximizing the clipped surrogate objective with a KL penalty to stabilize training:
 $J_{\mathrm{ChartSimRL}}(\theta)$ $=\mathbb{E}_{(I,T)\sim p_{\mathcal{D}},\,\{o_{i}\}_{i=1}^{M}\sim\pi_{\mathrm{old}}(\cdot|I,T)}\Bigg[\frac{1}{M}\sum_{i=1}^{M}\min\Big(\frac{\pi_{\theta}(o_{i}|I,T)}{\pi_{\mathrm{old}}(o_{i}|I,T)}\hat{A}_{i},$
@@
 ChartSimRL guides the model to generate chart code that better aligns with the original charts’ semantic and visual properties, significantly improving chart-to-code generation performance beyond what is achievable by supervised fine-tuning alone.
+Summary.
+ReChartPrompt and ChartSimRL have been effectively integrated into the ChartMaster framework.
+This framework not only leverages real-world data for enhanced data diversity but also employs a novel algorithm to ensure visual and semantic alignment in chart reproduction.
+Consequently, ChartMaster stands as a comprehensive solution for the chart-to-code generation task, demonstrating marked improvements in performance and generalization capabilities.
 4 Experiment
-4.1 Implementation Details
-During the collection of arXiv papers, we explicitly exclude papers that are used as benchmarks to avoid potential data leakage.
-In the data generation stage, we apply a greedy sampling strategy to filter chart data, retaining only images of 12 predefined chart types and discarding all others.
-Then, we randomly select an instruction from $\mathbf{P}_{\text{rechart}}$ to prompt the Qwen2.5-VL-72B bai2025qwen2 model to generate code via nucleus sampling, with a temperature of 0.1 and a top-p of 0.9.
-For training, we use the Qwen2.5-VL-7B model bai2025qwen2 in two stages.
-In Stage 1, we perform SFT on the entire ReChartPrompt-240K dataset with a learning rate of $2\times 10^{-5}$, batch size 128, and a cosine annealing scheduler for one epoch; the resulting model is saved for Stage 2.
-In Stage 2, ChartSimRL training is conducted on 10% of the ReChartPrompt-240K dataset, using a smaller learning rate of $5\times 10^{-6}$ and generating $M=4$ candidate codes per sample.
-Candidate sampling uses temperature 1.0, top-p 1.0, and top-k 80 to encourage diversity.
-The batch size remains 128 (32 samples $\times$ 4 candidates each).
-4.2 Evalution Benchmarks
-Following zhao2025chartcoder, we assess the model’s chart-to-code generation performance on multiple benchmarks.
-ChartMimic Direct Mimic Task yang2024chartmimic: This benchmark includes 600 chart images.
-GPT-4o scores (0–100) serve as high-level similarity metrics.
-Additionally, low-level F1 scores for text, layout, chart type, and color are computed from code execution for fine-grained analysis.
-Plot2Code Direct Asking wu2024plot2code: Metrics include code pass rate, text match rate, and a 10-point GPT-4V visual similarity score, jointly assessing code correctness and visual fidelity.
-ChartX Chart Redrawing Task xia2024chartx: This benchmark uses GPT-4 (0–5 scale) to evaluate code-generated chart redrawings.
 Table 1: Evaluation results of various MLLMs.
@@
 Closed-Source Model
-GeminiProVision team2023gemini 68.2 53.8 53.3 68.2 53.6 3.69 -
-Claude-3-opus anthropic2024claude 83.3 60.5 60.1 84.1 57.5 3.80 -
-GPT-4V hurst2024gpt 91.2 76.4 78.9 84.1 57.7 5.58 2.63
-GPT-4o hurst2024gpt 93.2 79.0 83.5 88.6 56.3 5.71 2.36
+GeminiProVision (Team et al., 2023) 68.2 53.8 53.3 68.2 53.6 3.69 -
+Claude-3-opus (Anthropic, 2024) 83.3 60.5 60.1 84.1 57.5 3.80 -
+GPT-4V (Hurst et al., 2024) 91.2 76.4 78.9 84.1 57.7 5.58 2.63
+GPT-4o (Hurst et al., 2024) 93.2 79.0 83.5 88.6 56.3 5.71 2.36
 Open-Source Model
-ChartAssisstant-13B meng2024chartassisstant - - - - - - 0.82
-ChartVLM-L-14B xia2024chartx 19.5 15.8 13.9 - - - 1.58
-DeepSeek-VL-7B lu2024deepseek 41.3 19.0 17.6 64.4 32.6 2.26 -
-TinyChart-3B zhang2024tinychart 42.5 26.3 25.9 43.2 44.6 2.19 1.89
-ChartLlama-13B han2023chartllama 57.5 24.8 28.1 58.4 40.3 2.32 0.94
-LLaVA-Next-Mistral-7B li2024llava 59.7 20.7 21.3 72.0 38.7 2.87 -
-InternVL2-8B chen2024internvl 61.8 34.4 38.9 77.3 37.1 2.78 1.63
-Qwen2-VL-7B wang2024qwen2 67.0 32.9 35.0 68.2 33.8 3.10 1.50
-MiniCPM-Llama3-V2.5-8B yao2024minicpm 80.3 36.6 42.1 76.3 37.3 2.61 1.66
-Qwen2-VL-72B wang2024qwen2 73.3 54.4 50.9 72.0 53.4 4.26 1.69
-InternVL2-Llama3-76B chen2024internvl 83.2 54.8 62.2 85.6 46.6 3.89 1.74
-Qwen2.5-VL-72B bai2025qwen2 88.5 72.7 71.8 84.8 68.4 6.83 2.52
-ChartCoder-7B zhao2025chartcoder 91.4 77.4 74.0 87.9 54.5 4.50 2.09
-Qwen2.5-VL-7B (Baseline) bai2025qwen2 65.5 39.9 34.2 67.4 43.8 4.60 2.18
-ChartMaster-7B 93.8 78.2 77.3 88.2 62.6 5.65 2.46
+ChartAssisstant-13B (Meng et al., 2024) - - - - - - 0.82
+ChartVLM-L-14B (Xia et al., 2024) 19.5 15.8 13.9 - - - 1.58
+DeepSeek-VL-7B (Lu et al., 2024) 41.3 19.0 17.6 64.4 32.6 2.26 -
+TinyChart-3B (Zhang et al., 2024b) 42.5 26.3 25.9 43.2 44.6 2.19 1.89
+ChartLlama-13B (Han et al., 2023) 57.5 24.8 28.1 58.4 40.3 2.32 0.94
+LLaVA-Next-Mistral-7B (Li et al., 2024a) 59.7 20.7 21.3 72.0 38.7 2.87 -
+InternVL2-8B (Chen et al., 2024) 61.8 34.4 38.9 77.3 37.1 2.78 1.63
+Qwen2-VL-7B (Wang et al., 2024a) 67.0 32.9 35.0 68.2 33.8 3.10 1.50
+MiniCPM-Llama3-V2.5-8B (Yao et al., 2024) 80.3 36.6 42.1 76.3 37.3 2.61 1.66
+Qwen2-VL-72B (Wang et al., 2024a) 73.3 54.4 50.9 72.0 53.4 4.26 1.69
+InternVL2-Llama3-76B (Chen et al., 2024) 83.2 54.8 62.2 85.6 46.6 3.89 1.74
+Qwen2.5-VL-72B (Bai et al., 2025) 88.5 72.7 79.1 84.8 68.4 6.83 2.52
+ChartCoder-7B (Zhao et al., 2025) 91.4 77.4 74.0 87.9 54.5 4.50 2.09
+Qwen2.5-VL-7B (Baseline) (Bai et al., 2025) 65.5 39.9 40.7 67.4 43.8 4.60 2.18
+ChartMaster-7B 93.8 78.2 85.1 88.2 62.6 5.65 2.46
 4.1 Comparison with SOTA
-Table 1 compares our ChartMaster-7B with various MLLMs on multiple chart-to-code benchmarks.
-ChartMaster-7B achieves SOTA results among open-source 7B-scale models, even approaching the performance of GPT-4o.
-Notably, ChartMaster-7B consistently outperforms the baseline Qwen2.5-VL-7B across all metrics; for example, on the ChartMimic benchmark, it improves both low-level and high-level metrics by nearly 40 percentage points, highlighting the effectiveness of our approach.
-Furthermore, although our training dataset is sourced from the much larger Qwen2.5-VL-72B model—essentially a distillation-like setting—ChartMaster-7B still surpasses Qwen2.5-VL-72B on several benchmarks.
-These results demonstrate the strength of the ChartMaster framework in narrowing, and in some cases closing, the performance gap with significantly larger models.
-Table 2: Ablation study on the contribution of each key component.
-ReChartPrompt ChartSimRL ChartMimic Plot2Code ChartX
-Exec.Rate Low-Level High-Level Pass Rate Text-Match Rating GPT-score
-65.5 39.9 34.2 67.4 43.8 4.60 2.18
-$\checkmark$ 91.1 73.7 73.3 80.3 59.3 5.34 2.36
-$\checkmark$ $\checkmark$ 93.8 78.2 77.3 88.2 62.6 5.65 2.46
+We instantiate ChartMaster on the Qwen2.5-VL-7B backbone, resulting in the ChartMaster-7B model, and conduct comprehensive comparisons with a range of MLLMs.
+The detailed implementation and evaluation protocols are provided in the Appendix A.
+As shown in Table 1, ChartMaster-7B achieves state-of-the-art performance among open-source models at the 7B scale, showing competitive performance against GPT-4o.
+Notably, ChartMaster-7B consistently outperforms the baseline Qwen2.5-VL-7B across all metrics; for instance, in the ChartMimic benchmark, it improves both low-level and high-level metrics by about 40 percentage points.
+Furthermore, although our training dataset is derived from the larger Qwen2.5-VL-72B model—essentially a distillation-like setting—ChartMaster-7B still surpasses Qwen2.5-VL-72B on several benchmarks.
+These results convincingly demonstrate the effectiveness of the ChartMaster framework.
 4.2 Ablation Study
@@
 The base Qwen2.5-VL-7B model, without ReChartPrompt or ChartSimRL, demonstrates limited performance across benchmarks, revealing its restricted ability in both code generation and visual/semantic understanding.
-Introducing the ReChartPrompt-240K dataset leads to significant improvements in all metrics, demonstrating the high quality and effectiveness of ReChartPrompt-240K for chart-to-code generation.
-Further applying ChartSimRL on top of ReChartPrompt brings consistent gains, especially in GPT-based evaluation scores.
-These results demonstrate that ChartSimRL effectively guides the model to produce code that better matches the original charts in both semantic and visual aspects.
+SFT with the ReChartPrompt-240K dataset leads to significant improvements in all metrics, demonstrating the high quality and effectiveness of ReChartPrompt-240K for chart-to-code generation.
+Further applying ChartSimRL on top of ReChartPrompt brings consistent gains, demonstrating its effectiveness in guiding the model to generate code that more accurately captures the original charts both semantically and visually.
+These findings confirm the complementary strengths of ReChartPrompt and ChartSimRL in enhancing ChartMaster’s overall performance.
+Table 2: Ablation study on the contribution of each key component.
+ReChartPrompt ChartSimRL ChartMimic Plot2Code ChartX
+Exec.Rate Low-Level High-Level Pass Rate Text-Match Rating GPT-score
+65.5 39.9 40.7 67.4 43.8 4.60 2.18
+$\checkmark$ 91.1 73.7 80.9 80.3 59.3 5.34 2.36
+$\checkmark$ $\checkmark$ 93.8 78.2 85.1 88.2 62.6 5.65 2.46
 Table 3: Ablation study of the Attribute and Visual similarity components in ChartSimRL.
@@
 Exec.Rate Low-Level High-Level
-91.1 73.7 73.3
-$\checkmark$ 92.1 76.2 75.2
-$\checkmark$ 92.1 77.7 76.0
-$\checkmark$ $\checkmark$ 93.8 78.2 77.3
+91.1 73.7 80.9
+$\checkmark$ 92.1 76.2 83.9
+$\checkmark$ 92.1 77.7 84.3
+$\checkmark$ $\checkmark$ 93.8 78.2 85.1
 Table 4: Ablation study of different attribute similarity metrics on the ChartMimic benchmark.
@@
 Exec.Rate Low-Level High-Level
-Baseline - 91.1 73.7 73.3
-Precision $\frac{|\mathcal{A}_{i}\cap\mathcal{A}^{*}|}{|\mathcal{A}_{i}|}$ 90.0 72.6 72.9
-Recall $\frac{|\mathcal{A}_{i}\cap\mathcal{A}^{*}|}{|\mathcal{A}^{*}|}$ 90.6 74.7 74.8
-F1 $\frac{|\mathcal{A}_{i}\cap\mathcal{A}^{*}|}{(|\mathcal{A}_{i}|+|\mathcal{A}^{*}|)/2}$ 91.6 75.4 75.3
-Jaccard $\frac{|\mathcal{A}_{i}\cap\mathcal{A}^{*}|}{|\mathcal{A}_{i}\cup\mathcal{A}^{*}|}$ 92.1 76.2 75.2
+- - 91.1 73.7 80.9
+Precision $\frac{|\mathcal{A}_{i}\cap\mathcal{A}^{*}|}{|\mathcal{A}_{i}|}$ 90.0 72.6 79.0
+Recall $\frac{|\mathcal{A}_{i}\cap\mathcal{A}^{*}|}{|\mathcal{A}^{*}|}$ 90.6 74.7 81.7
+F1 $\frac{|\mathcal{A}_{i}\cap\mathcal{A}^{*}|}{(|\mathcal{A}_{i}|+|\mathcal{A}^{*}|)/2}$ 91.6 75.4 84.5
+Jaccard $\frac{|\mathcal{A}_{i}\cap\mathcal{A}^{*}|}{|\mathcal{A}_{i}\cup\mathcal{A}^{*}|}$ 92.1 76.2 83.9
 Ablation study on ChartSimRL.
-ChartSimRL introduces a novel multimodal chart similarity reward, which goes beyond conventional text-based rewards by integrating both semantic and visual similarity between the candidate and original charts.
-To analyze the contribution of each component, we conduct ablation experiments as shown in Table 4.
-The results indicate that using either $R_{i}^{\mathrm{attr}}$ or $R_{i}^{\mathrm{vis}}$ alone improves all performance.
-Among them, the visual similarity reward leads to more pronounced gains, highlights the critical role of visual fidelity in chart-to-code generation.
-Furthermore, combining both rewards yields the best results, demonstrating the effectiveness of a multi-faceted reward design.
+ChartSimRL introduces a novel multimodal chart similarity reward that combines both semantic similarity ($R_{i}^{\mathrm{attr}}$) and visual similarity ($R_{i}^{\mathrm{vis}}$) between the candidate and original charts.
+To dissect the contribution of each component, we conduct ablation experiments summarized in Table 4.
+The results show that employing either $R_{i}^{\mathrm{attr}}$ or $R_{i}^{\mathrm{vis}}$ alone consistently improves performance across all evaluated metrics.
+Notably, the visual similarity reward yields more substantial gains, underscoring the critical importance of preserving visual fidelity in chart-to-code generation.
+Moreover, combining both rewards achieves the best overall results, demonstrating the advantage of a multi-faceted reward design that simultaneously captures semantic and visual aspects.
 Ablation study on Attribute Similarity.
-We further evaluate the impact of different attribute similarity metrics—Precision, Recall, F1 score, and Jaccard similarity—on the ChartMimic benchmark (Table 4).
-Results show that optimizing solely for Precision can slightly decrease certain metrics, indicating that focusing only on attribute correctness without considering coverage may be suboptimal.
-In contrast, Recall, F1, and Jaccard similarity all lead to better performance, with Jaccard similarity achieving the highest overall scores.
-Notably, while the F1 score balances Precision and Recall through the harmonic mean, Jaccard similarity measures the intersection over union between predicted and reference attribute sets, imposing a stricter penalty on missing or redundant attributes.
-This stricter overlap requirement allows Jaccard similarity to better capture the true semantic similarity between attribute sets, thus more effectively guiding the model’s learning process.
+We adopt Jaccard similarity as a stringent metric for attribute similarity, whereby a candidate table achieves a perfect score only if its attribute set exactly matches that of the ground truth; even minor discrepancies incur penalties.
+To thoroughly assess the impact of different attribute similarity measures—Precision, Recall, F1 score, and Jaccard similarity—we conduct experiments on the ChartMimic benchmark, with results summarized in Table 4.
+Our findings indicate that optimizing exclusively for Precision may lead to a slight decline in overall performance, as the model can achieve high Precision by predicting a limited subset of correct attributes while neglecting overall coverage.
+In contrast, Recall emphasizes coverage, which helps mitigate this issue and yields modest improvements.
+The F1 score, by harmoniously balancing Precision and Recall, further alleviates extreme biases and delivers enhanced overall performance.
+Notably, Jaccard similarity, measuring the intersection over union between predicted and reference attribute sets, enforces stricter penalties on both missing and redundant attributes.
+This higher overlap requirement enables Jaccard similarity to more faithfully capture the true semantic similarity between attribute sets, thereby resulting in the best overall performance.
 Ablation Study on Visual Similarity.
-Table 5 presents an ablation study of various visual similarity metrics on the ChartMimic benchmark.
-Standard metrics like MSE, SSIM wang2004image, and PSNR hore2010image, mainly assessing pixel-level or structural similarity, generally perform worse than advanced methods.
-In particular, SSIM underperforms significantly, indicating that pixel-based measures struggle to capture the complex visual nuances required for chart generation.
-In contrast, CNN-based metrics like AlexNet krizhevsky2012imagenet, VGG simonyan2014very and ResNet he2016deep that compare features in a learned representation space consistently outperform the baseline and pixel-level metrics across all evaluation criteria.
-ResNet-18 achieves the highest performance, highlighting the effectiveness of deep visual features.
+We use ResNet-18 (He et al., 2016) to extract features from charts to compute visual similarity.
+In fact, there are numerous methods to measure the similarity between two charts.
+To investigate the impact of different visual similarity metrics on model performance, we conduct an ablation study summarized in Table 5.
+Standard metrics such as MSE, SSIM (Wang et al., 2004), and PSNR (Hore & Ziou, 2010) primarily evaluate pixel-level or structural similarity (details in Appendix B).
+The table shows that these metrics generally perform worse than more advanced methods.
+Notably, SSIM exhibits a significant decline in performance, indicating that pixel-based measures struggle to capture the complex visual nuances necessary for effective chart-to-code generation.
+In contrast, CNN-based metrics like AlexNet (Krizhevsky et al., 2012), VGG (Simonyan & Zisserman, 2014), and ResNet (He et al., 2016), which compare features in a learned representation space, consistently outperform both the baseline and pixel-level metrics across all evaluation criteria.
+Among them, ResNet-18 achieves the highest performance, highlighting the effectiveness of deep visual features.
+For MLLM-based metrics, we leverage the high-level similarity prompt from ChartMimic combined with Qwen-2.5-VL-72B to evaluate the similarity between generated and reference charts.
+These metrics show improvements over the baseline.
+However, they still fall short of the best CNN-based metrics, suggesting that although MLLMs possess strong semantic understanding, further optimization is required for specialized visual tasks such as chart-to-code generation.
 Table 5: Ablation study of different visual similarity metrics on the ChartMimic benchmark.
@@
 Exec.Rate Low-Level High-Level
-Baseline 91.1 73.7 73.3
+- 91.1 73.7 80.9
 Standard Metrics:
-MSE 91.1 73.6 72.5
-SSIM 82.5 65.2 67.0
-PSNR 91.4 75.1 73.4
+MSE 91.1 73.6 77.9
+SSIM 82.5 65.2 74.6
+PSNR 91.4 75.1 82.1
 CNN-Based Metrics:
-AlexNet 90.3 74.7 74.1
-VGG 91.3 75.5 74.9
-ResNet-18 92.1 77.7 76.0
+AlexNet 90.3 74.7 82.6
+VGG 91.3 75.5 83.3
+ResNet-18 92.1 77.7 84.3
 MLLM-Based Metrics:
-Qwen-2.5-VL-72B 91.7 77.5 74.7
-For MLLM-based metrics, we use high-level similarity prompts from ChartMimic with Qwen-2.5-VL-72B to evaluate the similarity between generated and reference charts.
-These metrics further improve upon the baseline, especially in Execution Rate and Low-Level similarity.
-However, they still fall short of the best CNN-based metrics in High-Level similarity, suggesting that while MLLMs have strong semantic understanding, further optimization is needed for specialized visual tasks.
-Figure 3: Comparison of dataset diversity and model performance of Qwen2.5-VL-7B after fine-tuning with Chart2Code-160K and ReChartPromp-160K datasets.
+Qwen-2.5-VL-72B 91.7 77.5 83.9
+Figure 3: Comparison of diversity and fine-tuning results between Chart2Code-160K and ReChartPrompt-160K datasets.
 Comparison with Advanced Dataset.
-To thoroughly evaluate the diversity and quality of our dataset, we compare it with Chart2Code-160K.
-For fairness, 160K samples are randomly selected from our data to form the ReChartPromp-160K subset.
-Results are shown in Fig.
-3.
-The left subfigure displays normalized counts of unique attributes—Numeric, Color, Text, and Layout—extracted from both datasets using tool $\mathcal{G}(\cdot)$.
-Higher counts indicate greater attribute diversity.
-ReChartPromp-160K exhibits substantially more unique attributes across all categories, especially in Text and Layout.
-While Chart2Code-160K relies on seed data, often resulting in repetitive attributes (see Appendix Fig.
-5), ReChartPromp-160K draws from 30,071 diverse arXiv papers, yielding richer and less redundant attribute distributions.
-This greater diversity leads to tangible benefits: as shown in the right subfigure, models trained on ReChartPromp-160K consistently outperform those trained on Chart2Code-160K, demonstrating the importance of attribute diversity for robust and effective chart understanding and code generation.
+To comprehensively evaluate the diversity and quality of our dataset, we compare it with Chart2Code-160K (Zhao et al., 2025).
+For a fair comparison, we randomly sample 160K instances from our full dataset to construct the ReChartPromp-160K subset.
+Using the attribute extraction tool $\mathcal{G}(\cdot)$, we count unique chart attributes—including numerical values, colors, textual elements, and layouts—in both datasets.
+A higher number of unique attributes indicates greater attribute diversity.
+As shown in the left panel of Fig.
+3, ReChartPromp-160K exhibits a substantially richer attribute distribution across all categories, notably in text and layout.
+This advantage stems primarily from Chart2Code-160K’s reliance on seed data sources, which results in repeated attribute patterns, whereas ReChartPromp-160K samples from distinct arXiv papers, ensuring broader coverage and less redundancy (see Appendix Fig.
+5).
+This higher diversity brings clear benefits: as illustrated in the right panel, models fine-tuned on ReChartPromp-160K consistently outperform those trained on Chart2Code-160K, demonstrating the importance of attribute diversity for robust and effective chart understanding and code generation.
 4.3 Qualitative analysis
@@
 Incorporating ReChartPrompt significantly enhances the chart-to-code generation capability of the base model, while ChartSimRL further improves the handling of fine details.
-To further evaluate the effectiveness of our proposed approach, Fig.
-4 presents qualitative comparisons of chart outputs generated by different models on the ChartMimic benchmark.
-The results reveal several key insights: (1) The baseline model can generate basic chart layouts, but often fails to reproduce fine-grained visual details.
-Consequently, its outputs frequently diverge from the reference charts.
-(2) ReChartPrompt-240K enhances chart-to-code generation accuracy.
-Fine-tuning the base model with our ReChartPrompt-240K dataset (“Base.+ReCha.”) makes the generated charts more accurate and diverse.
-This shows that using real-world chart-code pairs helps the model better adapt to different chart types and styles.
-However, there are still minor issues, like slight color or element position mismatches, indicating that standard supervised learning alone cannot achieve perfect visual consistence.
-(3) ChartSimRL further improves visual and semantic alignment.
-Incorporating ChartSimRL algorithm enables the model to more closely align its outputs with the ground truth.
-For example, the color usage in chart 1 and the arrow style in chart 2 better match the ground truth.
-In summary, by combining ReChartPrompt and ChartSimRL, ChartMaster achieves visually accurate and semantically precise chart code, demonstrating strong generalization.
-This integrated approach also significantly narrows the gap with models like GPT-4o, highlighting ChartMaster’s practical effectiveness in chart-to-code generation.
+Based on extensive experiments, we observe that ReChartPrompt generates charts with diverse and rich attributes, enabling the construction of a high-quality dataset that substantially enhances model performance.
+Building upon the distinctive features of the chart-to-code generation task, we propose the ChartSimRL algorithm, which further enhances the model’s capabilities.
+To comprehensively analyze the improvements brought by these contributions, we conduct a qualitative comparison of generated charts at different training stages on the ChartMimic benchmark (Fig.
+4).
+Our key findings are summarized as follows: (1) The baseline model produces basic chart layouts but often fails to replicate fine-grained visual details, leading to noticeable discrepancies between generated outputs and reference charts.
+(2) Fine-tuning the base model on our ReChartPrompt-240K dataset (“Base.+ReCha.”) significantly improves chart-to-code generation accuracy.
+This improvement arises from the diverse, high-quality training data generated by conditioning on real-world chart prompts.
+Nonetheless, minor issues remain, such as slight mismatches in color or element positioning compared to the ground truth, indicating that supervised fine-tuning alone does not achieve perfect visual consistency.
+(3) Incorporating the ChartSimRL algorithm further improves both visual and semantic alignment.
+Notably, the model demonstrates enhanced color accuracy (as seen in the first column of Fig.
+4) and more faithful reproduction of arrow styles in the second column, reflecting improved attention to key factual details.
+(4) ChartMaster competes favorably with GPT-4o.
+Notably, the ChartMaster-7B model can generate charts that more closely resemble the ground truth than those from GPT-4o, especially excelling in “mimicking” chart attributes.
+Additional generation results in Appendix Fig.
+8 consistently support these conclusions.
 5 Conclusion
-In this paper, we propose ChartMaster, including a novel chart-to-code generation framework and a customized reinforcement learning algorithm.
-By introducing ReChartPrompt, we overcome the data homogeneity issues of previous works and construct a highly diverse ReChartPrompt-240K dataset.
-Our ChartSimRL algorithm integrates both semantic and visual similarity rewards, enabling the model to generate chart code that closely matches the original charts.
-Experimental results demonstrate that ChartMaster achieves performance competitive with GPT-4o on the chart-to-code generation task.
-Beyond its technical contributions, ChartMaster can facilitate automated scientific reporting and support data-driven decision-making across a wide range of domains.
-While our current framework focuses on common chart types and Python-based code, extending its capabilities to support a broader variety of chart formats and programming languages remains an important direction for future work.
-Appendix A Appendix
+In this paper, we propose ChartMaster, a novel chart-to-code generation framework paired with a tailored reinforcement learning algorithm.
+By introducing ReChartPrompt, we address data homogeneity issues in prior work and build a highly diverse ReChartPrompt-240K dataset.
+Our ChartSimRL algorithm combines semantic and visual similarity rewards, enabling the model to generate chart code that closely matches original visuals.
+Experiments show ChartMaster achieves performance on par with GPT-4o in chart-to-code tasks.
+We will open source all resources to foster community development and advance research in this area.
+Beyond its technical innovations, ChartMaster supports automated scientific reporting and empowers data-driven decision-making across a wide range of domains.
+While our current framework targets common chart types and Python-based code, expanding its scope to include a wider range of chart formats and programming languages is an exciting direction for future work.
+Appendix A Implementation and Evaluation Details
+During the collection of arXiv papers, we explicitly exclude papers that are used as benchmarks to avoid potential data leakage.
+In the data generation stage, we apply a greedy sampling strategy to filter chart data, retaining only images of 12 predefined chart types and discarding all others.
+Then, we randomly select an instruction from $\mathbf{P}_{\text{rechart}}$ to prompt the Qwen2.5-VL-72B (Bai et al., 2025) model to generate code via nucleus sampling, with a temperature of 0.1 and a top-p of 0.9.
+For training, we use the Qwen2.5-VL-7B model (Bai et al., 2025) in two stages.
+In Stage 1, we perform SFT on the entire ReChartPrompt-240K dataset with a learning rate of $2\times 10^{-5}$, batch size 128, and a cosine annealing scheduler for one epoch; the resulting model is saved for Stage 2.
+In Stage 2, ChartSimRL training is conducted on 10% of the ReChartPrompt-240K dataset, using a smaller learning rate of $5\times 10^{-6}$ and generating $M=4$ candidate codes per sample.
+Candidate sampling uses temperature 1.0, top-p 1.0, and top-k 80 to encourage diversity.
+The batch size remains 128 (32 samples $\times$ 4 candidates each).
+For evalution, we assess the model’s chart-to-code generation performance on multiple benchmarks.
+ChartMimic Direct Mimic Task (Yang et al., 2024a): This benchmark includes 600 chart images.
+GPT-4o scores (0–100) serve as high-level similarity metrics.
+Additionally, low-level F1 scores for text, layout, chart type, and color are computed from code execution for fine-grained analysis.
+Plot2Code Direct Asking (Wu et al., 2024): Metrics include code pass rate, text match rate, and a 10-point GPT-4V visual similarity score, jointly assessing code correctness and visual fidelity.
+ChartX Chart Redrawing Task (Xia et al., 2024): This benchmark uses GPT-4 (0–5 scale) to evaluate code-generated chart redrawings.
+Appendix B Standard Metrics
+We consider two RGB images: the original chart image $I_{i}\in\mathbb{R}^{H\times W\times 3}$ and the generated chart image $\hat{I}_{i}\in\mathbb{R}^{H\times W\times 3}$, where $H$ and $W$ denote the height and width of the images respectively (both images are resized to the same height and width before comparison), and $3$ corresponds to the RGB color channels.
+Below, we describe how to quantify the visual similarity between $I_{i}$ and $\hat{I}_{i}$ using metrics such as Mean Squared Error (MSE), Structural Similarity (SSIM) (Wang et al., 2004), and Peak Signal-to-Noise Ratio (PSNR) (Hore & Ziou, 2010).
+B.1 Mean Squared Error
+The Mean Squared Error (MSE) is defined as:
+$\mathrm{MSE}(I_{i},\hat{I}_{i})=\frac{1}{H\times W\times 3}\sum_{h=1}^{H}\sum_{w=1}^{W}\sum_{c=1}^{3}\left(I_{i}(h,w,c)-\hat{I}_{i}(h,w,c)\right)^{2}$
+This formula computes the average squared difference between the pixel values of the two images over all spatial locations and color channels.
+A smaller MSE indicates higher similarity between $I_{i}$ and $\hat{I}_{i}$.
+To convert the MSE into a similarity score, we define the MSE-based similarity as:
+$\mathrm{MSE\_Similarity}=\frac{1}{1+\mathrm{MSE}(I_{i},\hat{I}_{i})}\in(0,1]$
+  • When $\mathrm{MSE}\to 0$, $\mathrm{MSE\_Similarity}\to 1$, indicating the images are almost identical.
+  • When $\mathrm{MSE}\to\infty$, $\mathrm{MSE\_Similarity}\to 0$, indicating large differences between the images.
+B.2 Structural Similarity
+The Structural Similarity (SSIM) is a perceptual metric that quantifies the similarity between two images by comparing local patterns of pixel intensities.
+It is computed on local sliding windows centered at each pixel location.
+For each window, local statistics including mean, variance, and covariance are calculated to evaluate the similarity.
+The final SSIM value for each channel is obtained by averaging these local SSIM values over all spatial positions, and the overall SSIM between two RGB images is computed by averaging over the three color channels.
+Formally, for each color channel $c\in\{R,G,B\}$, the SSIM is defined as:
+$\mathrm{SSIM}_{c}(I_{i}^{c},\hat{I}_{i}^{c})=\frac{(2\mu_{I_{i}^{c}}\mu_{\hat{I}_{i}^{c}}+C_{1})(2\sigma_{I_{i}^{c}\hat{I}_{i}^{c}}+C_{2})}{(\mu_{I_{i}^{c}}^{2}+\mu_{\hat{I}_{i}^{c}}^{2}+C_{1})(\sigma_{I_{i}^{c}}^{2}+\sigma_{\hat{I}_{i}^{c}}^{2}+C_{2})}$
+where
+  • $\mu_{I_{i}^{c}}$ and $\mu_{\hat{I}_{i}^{c}}$ are the local means computed within the sliding window.
+  • $\sigma_{I_{i}^{c}}^{2}$ and $\sigma_{\hat{I}_{i}^{c}}^{2}$ are the local variances.
+  • $\sigma_{I_{i}^{c}\hat{I}_{i}^{c}}$ is the local covariance.
+  • $C_{1}=(K_{1}L)^{2}$ and $C_{2}=(K_{2}L)^{2}$ are constants to stabilize the division, with default values $K_{1}=0.01$, $K_{2}=0.03$.
+$L$ is the dynamic range of the pixel values.
+For 8-bit grayscale images, $L=255$.
+In our implementation, all images are converted to np.float32 and normalized by dividing by 255, so the pixel values are in the range $[0,1]$.
+Therefore, $L=1.0$ is used for SSIM calculation.
+The overall mean SSIM between the two RGB images is then calculated by averaging over all spatial positions $(x,y)$ in each channel and then over the three channels:
+$\mathrm{SSIM}(I_{i},\hat{I}_{i})=\frac{1}{3}\sum_{c=1}^{3}\frac{1}{H\times W}\sum_{x=1}^{H}\sum_{y=1}^{W}\mathrm{SSIM}_{c}(I_{i}^{c}(x,y),\hat{I}_{i}^{c}(x,y))\in[0,1]$
+  • When $\mathrm{SSIM}\to 1$, the images are structurally almost identical.
+  • When $\mathrm{SSIM}\to 0$, there are significant structural differences between the images.
+B.3 Peak Signal-to-Noise Ratio
+Peak Signal-to-Noise Ratio (PSNR) is a widely used metric to measure the quality of reconstructed images compared to the original images. It is defined as:
+$\mathrm{PSNR}(I_{i},\hat{I}_{i})=10\log_{10}\left(\frac{L^{2}}{\mathrm{MSE}(I_{i},\hat{I}_{i})}\right)$
+where $L$ is the dynamic range of the pixel values. For normalized images in $[0,1]$, $L=1.0$.
+In practical scenarios, PSNR values typically range in tens of decibels and can vary widely, which may cause instability during optimization.
+To mitigate this effect, we normalize the PSNR values within each rollout batch by dividing them by the maximum PSNR value in that batch:
+$\mathrm{PSNR}_{\mathrm{norm}}(I_{i},\hat{I}_{i})=\frac{\mathrm{PSNR}(I_{i},\hat{I}_{i})}{\max_{\hat{I}_{j}\in\text{rollout batch}}\mathrm{PSNR}(I_{j},\hat{I}_{j})}\in(0,1]$
+  • When $\mathrm{PSNR}_{\mathrm{norm}}\to 1$, the reconstructed image $\hat{I}_{i}$ is very similar to the original image $I_{i}$.
+  • When $\mathrm{PSNR}_{\mathrm{norm}}\to 0$, there exist significant differences between the images.
 Figure 5: Dataset visualization.